QUESTION: SOLVE EACH OF THE FOLLOWING BY USING EITHER ADDITION OF SUBSTITUTION. IF A UNIQUE SOLUTION DOES NO EXIST, STATE WHETHER TH SYSTEM IS DEPENDENT OR CONSISTENT.
38. 10X+2Y=7
Y= -5X+3

Answer by funmath(2933)

QUESTION: SOLVE EACH OF THE FOLLOWING BY USING EITHER ADDITION Or SUBSTITUTION. IF A UNIQUE SOLUTION DOES NO EXIST, STATE WHETHER TH SYSTEM IS DEPENDENT OR inCONSISTENT.
38. 10X+2Y=7
Y= -5X+3
Since the 2nd equation is solved for y, lets use substitution. Substitute -5x+3 in for Y in the first equation and solve for x:
:
10X+2(-5X+3)=7
10X-10X+6=7
6=7
No solution.
Since you lost your variables and 6 does not = 7, that means that the system is inconsistent. Graphically, they are parallel lines that never intersect.
Happy Calculating!!!